OpenStudy (girlgoyleh.):

WILL FAN AND MEDAL! Which of the following expressions are equivalent to the expression below? 8x+32y

OpenStudy (anonymous):

is there a list of equations to choose from?

OpenStudy (anonymous):

divide by 8 .. then see if that matches one of your answers 8/8x +32/8y ... x + 4y

OpenStudy (girlgoyleh.):

4x+4y+16x+16y is an answer choice

OpenStudy (girlgoyleh.):

4(4x+8y)-2(4x)

OpenStudy (anonymous):

4x + 4y +16x+16y = 20x + 20y = x + y .. so, no 4(4x+8y) - 2(4x) = 16x +32y -8x = 8x +32y this is identical to your first equation, so yes